The tax rates for a particular year are shown below: Taxable Income Tax Rate $0 – 50,000 15 % 50,001 – 75,000 25 % 75,001 – 100,000 34 % 100,001 – 335,000 39 % What is the average tax rate for a firm with taxable income of $124,513?
Answer #1

Total taxes=(50,000*0.15)+(25,000*0.25)+(25,000*0.34)+(24,513*0.39)

= 31810.07

Average tax rate=Total taxes/Taxable income

= 31810.07/124,513

=25.55%(Approx).

NOTE:The first $50,000 is taxed at 15%;the next (75,000-50,000)=$25,000 at 25% and so on.